A block of mass M = 10 kg is subjected to a force
Answer #1

a)

here

Ffriction = m * g * uk

Ffriction = 10 * 9.8 * 0.25 = 24.5 N

then the force that give motion to the block is = 30 - 24.5 = 5.5 N

then the acceleation of the block is

5.5 = m * a

a = 5.5 / 10 = 0.55 m/s^2

b)

here also we use the same formula

Ffriciton = 24.5 N

F = 90 - 24.5 = 65.5

then the acceleation is

a = 65.5 / 10 = 6.55 m /s^2